# Leveraging Passage Embeddings for Efficient Listwise Reranking with Large Language Models

### v1

@@ Abstract
Recent studies have demonstrated the effectiveness of using large language language models (LLMs) in passage ranking. The listwise approaches, such as RankGPT, have become new state-of-the-art in this task. However, the efficiency of RankGPT models is limited by the maximum context length and relatively high latency of LLM inference. To address these issues, in this paper, we propose PE-Rank, leveraging the single passage embedding as a good context compression for efficient listwise passage reranking. By treating each passage as a special token, we can directly input passage embeddings into LLMs, thereby reducing input length. Additionally, we introduce an inference method that dynamically constrains the decoding space to these special tokens, accelerating the decoding process. For adapting the model to reranking, we employ listwise learning to rank loss for training. Evaluation results on multiple benchmarks demonstrate that PE-Rank significantly improves efficiency in both prefilling and decoding, while maintaining competitive ranking effectiveness. The Code is available at https://github.com/liuqi6777/pe_rank.
Leveraging Passage Embeddings for Efficient Listwise Reranking
with Large Language Models
Qi Liu1,2, Bo Wang2, Nan Wang2and Jiaxin Mao1 1Gaoling School of Artificial Intelligence, Renmin University of China, 2Jina AI {qiliu6777, maojiaxin}@gmail.com, {bo.wang, nan.wang}@jina.ai
1 Introduction
@@
Passage ranking, which aims to rank each passage in a large corpus according to its relevance to the user’s information need expressed in a short query, is an important task in IR and NLP and plays a crucial role in many applications such as web search and retrieval-augmented generation. To achieve both effectiveness and efficiency, current mainstream approaches usually follow a two-stage paradigm known as “retrieval-then-rerank”, which involves efficiently retrieving a set of candidates first, and further reranking them with a reranker to boost the effectiveness Nogueira et al. (2019).
@@
In the first retrieval stage, dense retrieval models based on a bi-encoder architecture are widely used Karpukhin et al. (2020). Trained on large-scale text pairs through contrastive learning, these models can encode text information into a low-dimensional dense embedding and capture semantic relevance using vector similarity.
@@
Figure 1: Comparison between RankGPT (upper) and PE-Rank (lower). RankGPT takes the whole passages as input and outputs ordered numbers, while PE-Rank takes a list of special tokens as both input and output. On the right side, we show the reranking results on DL19 using different forms of inputs.
@@
In the second reranking stage, we can employ more sophisticated models for better ranking performance. A common reranking model is a supervised model based on the cross-encoder design Nogueira et al. (2019). With the emergence of LLMs, such as GPT-4 OpenAI (2024), a series of studies have tried to leverage LLMs’ text comprehension and reasoning abilities for zero-shot reranking. Typically, there are three main prompting approaches: pointwise Liang et al. (2022); Sachan et al. (2022), pairwise Qin et al. (2023), and listwise Sun et al. (2023); Pradeep et al. (2023a). Among these methods, listwise approaches like RankGPT Sun et al. (2023) are regarded as the most effective, achieving state-of-the-art performance by directly producing a final ranking list for multiple passages, rather than merely assessing the relevance of a single passage or the relative position between two passages.
@@
While the listwise approaches demonstrate potential in the reranking task, they are limited by two challenges. Firstly, LLMs are limited by context length and cannot rank multiple passages simultaneously, necessitating techniques such as a sliding window strategy to complete the ranking process Sun et al. (2023). Secondly, incorporating entire passages into prompts significantly increases inference costs, resulting in high latency, which is untenable in the ranking scenario.
@@
To tackle these issues, it is imperative to compress listwise reranking prompts. Some context compression methods have been proposed for LLMs and can be categorized into two types: compressing the context into dense memory slots Mu et al. (2024); Chevalier et al. (2023); Ge et al. (2023) and directly editing the input contexts Jiang et al. (2023b). Nonetheless, existing methods exhibit relatively low compression rates and usually only compress a single passage, rendering them inadequate for ranking tasks.
@@
For, we first highlight that in the “retrieval-then-rerank” pipeline, dense retrieval models have been trained as effective text compressors with their embedding capable of representing nearly as much information as the original text Morris et al. (2023). From this perspective, in the paper, we propose a novel and efficient listwise passage reranking method named PE-Rank, leveraging the single embedding of the passage as the compressed representation. Specifically, we obtain the passage embedding from a dense retrieval model and regard it as a special token of the LLM to replace the original text as input. To align the embedding space of the retrieval model and the input embedding space of the LLM, we use a projector as a bridge between the two models, which is inspired by previous work about modality alignment Liu et al. (2024).
@@
To adapt PE-Rank to ranking tasks, we propose novel inference and training methods. For accurate and efficient inference, we propose a “Dynamic-Constrained Decoding” strategy that dynamically changes the decoding spaces to a set of special tokens that represent the rest of the passages to be ranked. We employ two-stage training, first training the projector for modality alignment, then training both the projector and LLM for ranking tasks using listwise learning to rank loss.
@@
We evaluate PE-Rank on popular retrieval benchmarks TREC DL and BEIR. Experimental results demonstrate that PE-Rank achieves comparable ranking performance to uncompressed methods while significantly improving inference efficiency. Notably, when reranking top 100 candidates retrieval by BM25 on DL19, NDCG@10 of PE-Rank is only reduced by less than 2% compared to the uncompressed method under the same settings while reducing the latency by a factor of 4.5.
@@
  • We propose a novel efficient listwise reranking method, PE-Rank, first using passage embeddings for context compression in ranking.
@@
Figure 2: Overview of PE-Rank under a two-stage ranking paradigm. (a) is retrieval stage, retrieve $n$ passage embeddings; (b) is the forward pass procedure of LLM; (c) shows the listwise decoding process.
2 Methodology
2.1 Overview
@@
Specifically, we first use the dense retrieval model to pre-encode the corpus into a vector index. Given a query $q$, we use the same encoder to encode it into an embedding and retrieve several most relevant candidate passages $\mathcal{P}_{cand}=[p_{1},...,p_{n}]$ and their embeddings $\bm{e}_{p_{1}},...,\bm{e}_{p_{n}}$. Here vector similarity is used as the relevance score between query and passages.
In the reranking stage, our key idea is to take the embeddings from the previous stage as a good context compression of passages. Therefore, we propose replacing the original passage with the single embedding representation as the input of LLMs. However, there are dimensional and distribution differences between the passage embeddings and LLM’s token embeddings, which require us to bridge the gap between two spaces with a learned mapping function. Taking inspiration from previous work on aligning two modalities Liu et al. (2024), we introduce a two-layer MLP, denoted as $\mathbf{E}_{M}$, as the mapping function. Here we treat these transformed embeddings $\mathbf{E}_{M}(\bm{e}_{p_{i}})$ as the embeddings of additional out-of-vocabulary special tokens, where one passage is represented as one special token, for example <p1> represents $p_{1}$.
@@
$\mathbf{E}_{\text{In}}^{(1)}=\mathbf{E}_{t}(I\oplus q)\oplus\mathbf{E}_{M}(\bm{e}_{p_{1}})\cdots\oplus\mathbf{E}_{M}(\bm{e}_{p_{n}}),$ (1)
where $\mathbf{E}_{t}$ is the token embedding layer of LLM. The complete prompts are listed in Appendix F. In the next section, we will introduce how to output the ranking list in detail.
@@
2.2 Inference
During inference, listwise rerankers aim to output a ranking list directly. For LLM-based listwise approaches, we usually generate the ranking list autoregressively. In previous work, LLMs are prompted to generate a string that could be parsed into a ranking list, such as “[2] > [3] > [1]…” Sun et al. (2023); Pradeep et al. (2023a). However, in early experiments, we found that generating a string representing ranking may be difficult and slow, as LLM may output in the wrong format or output useless content, such as explanation.
Input : Candidates $\mathcal{P}_{cand}=[p_{1},...,p_{n}]$, Intial Input Embeddings $\mathbf{E}_{\text{In}}^{(1)}$
@@ 3       $\bm{h}_{i}\leftarrow\text{Transformer}(\mathbf{E}_{\text{In}}^{(i)})$
Algorithm 1 DC Decoding
To address this issue, we propose a “Dynamic-Constrained Decoding” (DC Decoding for short) strategy in Algorithm 1. During decoding, we dynamically change the decoding spaces according to the rest of the passages that need to be ranked, treating the embedding representation of those passages as a special set of tokens. At each generation step, we no longer output a normal numerical token but instead constrain the decoding space only in these special tokens, to perform accurate ranking. Therefore, we can directly output a list of tokens that represent the ranking of passages, such as “<p2><p3><p1>…”. Furthermore, as the decoding space and the number of generated tokens are much smaller than the original vocabulary space, inference will be accelerated.
@@ 2.3 Training
$\mathbf{E}_{\text{In-Align}}=\mathbf{E}_{t}(I)\oplus\mathbf{E}_{M}(\bm{e}_{t}),$ (2)
@@
$\mathcal{L}_{\text{Align}}=-\sum_{i=1}\log P_{\theta}(t_{i}|\mathbf{E}_{\text{In-Align}}\oplus\mathbf{E}_{t}(t_{<i})).$ (3)
@@
Previous listwise approaches employed supervised fine-tuning (SFT) paradigms for training Pradeep et al. (2023a, b). By distilling from existing reranking models, the LLMs acquire ranking knowledge. However, the dynamic nature of the decoding space renders standard SFT inapplicable in this context.
Therefore, we propose that the decoding process can be viewed as a sequential ranking learning process: at each step, we provide the previously decoded rankings and maximize the probability of generating the next most relevant passage. Formally, if given a query $q$ and the golden ranking list $[p_{1},...,p_{n}]$, at step $i$, we maximize the conditional probability of $p_{i}$ given $q$ and previous $p_{<i}$:
$P_{\theta}(p_{i}|q,p_{<i})$ $=P_{\theta}(p_{i}|\mathbf{E}_{\text{In}}^{(i)})$ (4)
@@
where $\theta$ is the model’s parameters. Considering the whole sequential process, it is equivalent to listwise learning to rank loss ListMLE Xia et al. (2008):
$\mathcal{L}_{\text{rank}}=-\sum_{i=1}^{n}\log P_{\theta}(p_{i}|\mathbf{E}_{\text{In}}^{(i)}).$ (5)
Here we only leverage the passage embeddings for ranking, as illustrated in the prompt (a) in Figure 3. The full prompts can be found in Appendix F.
Figure 3: Illustration of two types of training data and the learning-to-rank training process.
@@
$\mathcal{L}_{\text{content}}=-\sum_{i=1}^{n}\log P_{\theta}(p_{i}|\mathbf{E}_{\text{In-c}}^{(i)}),$ (6)
@@
$\mathcal{L}_{\text{KL}}=\sum_{i=1}^{n}{D}_{\text{KL}}(P_{\theta}(p_{i}|\mathbf{E}_{\text{In}}^{(i)})\|P_{\theta}(p_{i}|\mathbf{E}_{\text{In-c}}^{(i)}).$ (7)
@@
$\mathcal{L}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{content}}+\alpha\mathcal{L}_{\text{KL}}.$ (8)
Here $\alpha$ is set to 0.2. We fine-tune both MLP and LLM in this stage but remain encoder frozen.
@@
3 Experiment Setup
3.1 Evaluation Datasets
We evaluate PE-Rank on multiple retrieval benchmarks, including TREC DL Craswell et al. (2020) and BEIR Thakur et al. (2021). TREC DL uses the MS MARCO dataset Bajaj et al. (2016) as the retrieval corpus and has fine-grained relevance annotations. We use the test sets of TREC DL 2019 and TREC DL 2020, which contain 43 and 54 queries respectively. BEIR contains 18 datasets from different fields with different query requirements, aiming to evaluate the generalization ability of ranking models. Following previous work Sun et al. (2023), we conduct evaluations on 8 datasets that contain a relatively small number of queries. We use nDCG@10 as evaluation metrics.
3.2 Implementation Details

### v2

@@ Abstract.
Recent studies have demonstrated the effectiveness of using large language language models (LLMs) in passage ranking. The listwise approaches, such as RankGPT, have become new state-of-the-art in this task. However, the efficiency of RankGPT models is limited by the maximum context length and relatively high latency of LLM inference. To address these issues, in this paper, we propose PE-Rank, leveraging the single passage embedding as a good context compression for efficient listwise passage reranking. By treating each passage as a special token, we can directly input passage embeddings into LLMs, thereby reducing input length. Additionally, we introduce an inference method that dynamically constrains the decoding space to these special tokens, accelerating the decoding process. For adapting the model to reranking, we employ listwise learning to rank loss for training. Evaluation results on multiple benchmarks demonstrate that PE-Rank significantly improves efficiency in both prefilling and decoding, while maintaining competitive ranking effectiveness. [^1] The code is available at https://github.com/liuqi6777/pe_rank.
Reranking, Large Language Models, Efficiency
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714554isbn: 979-8-4007-1274-6/25/04ccs: Information systems Language models
1. Introduction
@@
Passage ranking, which aims to rank each passage in a large corpus according to its relevance to the user’s information need expressed in a short query, is an important task in information retrieval and natural language processing and plays a crucial role in many applications such as web search and retrieval-augmented generation. To achieve both effectiveness and efficiency, current mainstream approaches usually follow a two-stage paradigm known as “retrieval-then-rerank”, which involves efficiently retrieving a set of candidates first, and further reranking them with a reranker to boost the effectiveness (Matveeva et al., 2006; Nogueira et al., 2019).
@@
In the first retrieval stage, dense retrieval models based on a bi-encoder architecture are widely used (Karpukhin et al., 2020). Trained on large-scale datasets of text pairs through contrastive learning, these models can encode text into a low-dimensional dense embedding and capture the relevance between query and passage using vector similarity.
@@
Figure 1. Comparison between RankGPT (upper) and PE-Rank (lower). RankGPT takes the whole passage as input and outputs ordered numbers, while PE-Rank takes a list of special tokens as both input and output. On the right side, we show the reranking results on DL19 using different forms of inputs.
@@
In the second reranking stage, we can employ more sophisticated models for better ranking performance. A common reranking model is a supervised model based on the cross-encoder design (Nogueira et al., 2019). With the emergence of large language models (LLMs), such as GPT-4 (OpenAI, 2024), a series of studies have tried to leverage LLMs’ text comprehension and reasoning abilities for zero-shot reranking. Typically, there are three main prompting approaches: pointwise (Liang et al., 2022; Sachan et al., 2022), pairwise (Qin et al., 2023), and listwise (Sun et al., 2023; Pradeep et al., 2023a). Among these methods, listwise approaches like RankGPT (Sun et al., 2023) have achieved state-of-the-art performance by directly producing a final ranking list for multiple passages, rather than merely assessing the relevance of a single passage or the relative position between two passages.
@@
While the listwise approaches demonstrate good performance in the reranking task, they are limited by two challenges. Firstly, some LLMs are limited by context length and cannot rank multiple passages simultaneously, necessitating techniques such as a sliding window strategy to complete the ranking process (Sun et al., 2023). Secondly, incorporating entire passages into prompts significantly increases inference costs, resulting in high latency in practice (Chen et al., 2024), which is untenable in the ranking scenario.
@@
To tackle these issues, it is imperative to compress listwise reranking prompts. Some context compression methods have been proposed for LLMs and can be categorized into two types: compressing the context into dense memory slots (Mu et al., 2024; Chevalier et al., 2023; Ge et al., 2023) and directly editing the input contexts (Jiang et al., 2023b). Nonetheless, existing methods exhibit relatively low compression rates and usually only compress a single passage, rendering them inadequate for ranking tasks.
@@
For compressing multiple passages for reranking, we first highlight that in the “retrieval-then-rerank” pipeline, dense retrieval models have been trained as effective text compressors with their embedding capable of representing nearly as much information as the original text (Morris et al., 2023). In this paper, we propose a novel and efficient listwise passage reranking method named PE-Rank, leveraging the single embedding of the passage as the compressed representation. Specifically, we obtain the passage embedding from a dense retrieval model and regard it as a special token of the LLM to replace the original text as input. To align the embedding space of the retrieval model and the input embedding space of the LLM, we use a projector as a bridge between the two models, which is inspired by previous work about modality alignment (Liu et al., 2024a).
@@
To enable PE-Rank to complete ranking tasks, we propose novel inference and training methods. For efficient inference, we propose a “Dynamic-Constrained Decoding” strategy that dynamically changes and constrains the decoding spaces to a set of special tokens that represent the rest of the passages to be ranked. We employ two-stage training, first training the projector for modality alignment, then training both the projector and LLM for ranking tasks using listwise learning to rank loss.
@@
We evaluate PE-Rank on popular retrieval benchmarks TREC DL and BEIR. Experimental results demonstrate that PE-Rank achieves comparable ranking performance to uncompressed methods while notably improving inference efficiency. Notably, when reranking top 100 candidates retrieval by BM25 on DL19, NDCG@10 of PE-Rank is only reduced by less than 2% compared to the uncompressed method under the same settings while reducing the latency by a factor of 4.5.
@@
  • We propose a novel efficient listwise reranking method, PE-Rank, which is the first model that leverages passage embeddings for context compression and highly efficient listwise reranking.
  • We propose a two-stage training strategy that includes alignment and learning-to-rank stage to effectively train PE-Rank and a novel decoding method for efficient inference.
@@
Figure 2. Overview of PE-Rank under a two-stage ranking paradigm. (a) is retrieval stage, retrieve $n$ passage embeddings; (b) is the forward pass procedure of LLM; (c) shows the listwise decoding process.
2. Related Work
2.1. Large Language Models as Rerankers
Recently, large language models have demonstrated impressive effectiveness on many tasks. Many studies also attempt to utilize LLMs for zero-shot reranking. In general, there are three paradigms for prompting LLMs: pointwise, pairwise, and listwise.
The pointwise approach evaluates the relevance score on one query-passage pair at a time, including relevance generation (Liang et al., 2022; Liu et al., 2024c) and query generation (Sachan et al., 2022). The pairwise approach prompts LLM with a pair of passages to a given query to indicate which is more relevant, using aggregation methods (Pradeep et al., 2021) or sorting algorithms (Qin et al., 2023; Zhuang et al., 2023; Yoon et al., 2024) to derive the final ranking. The listwise approach aims to receive a query along with a list of candidates and directly generate a ranking list based on their relevance to the query (Ma et al., 2023; Sun et al., 2023). Recently, some studies have attempted to distill smaller listwise reranking models from existing powerful rerankers like RankGPT (Pradeep et al., 2023a, b; Zhang et al., 2023; Liu et al., 2024b). Our proposed method aims to enhance the efficiency of listwise approaches while preserving their effectiveness.
2.2. Context Compression
Context compression, which seeks to reduce the input length of LLMs while retaining the essential information from the original context, has recently garnered considerable attention. One approach is to heuristic modify the context to make it concise while retaining key information. LLMLingua (Jiang et al., 2023b) introduces a coarse-to-fine prompt compression method based on the perplexity score. RECOMP (Xu et al., 2023) proposes compressing documents into text summaries for RAG. Another direction is to compress the text into dense slots or soft prompts, such as AutoCompressor (Chevalier et al., 2023), ICAE (Ge et al., 2023), and Gist (Mu et al., 2024). However, these methods only compress a single prompt and are inadequate for ranking tasks. In contrast, our proposed method is specifically designed for ranking tasks and can be regarded as a variant of the second kind of method.
Recently, a contemporary work, xRAG, proposed using embedding models to compress a document into a token for RAG, which is similar to our proposed method (Cheng et al., 2024). Compared to it, our proposed PE-Rank method has the following differences: firstly, we compress prompts for the ranking task which is more complex, and secondly, we compress multiple documents as input at once.
3. Methodology
3.1. Overview
@@
Specifically, we first use the dense retrieval model to pre-encode the corpus into a vector index. Given a query $q$, we use the same encoder to encode it into an embedding and retrieve several most relevant candidate passages $\mathcal{P}_{cand}=[p_{1},...,p_{n}]$ and their embeddings $\bm{e}_{p_{1}},...,\bm{e}_{p_{n}}$. Depending on the retrieval model, the embedding of [CLS] token or the mean pooling is used. Then vector similarity is used as the relevance score between query and passages.
In the reranking stage, our key idea is to take the embeddings from the previous stage as a good context compression of passages. Therefore, we propose replacing the original passage with the single embedding representation as the input of LLMs. However, there are dimensional and distribution differences between the passage embeddings and LLM’s token embeddings, which require us to bridge the gap between two spaces with a learned mapping function. Taking inspiration from previous work on aligning two modalities (Liu et al., 2024a), we introduce a two-layer multi-layer perception (MLP), denoted as $\mathbf{E}_{M}$, as the mapping function. Here we treat these transformed embeddings $\mathbf{E}_{M}(\bm{e}_{p_{i}})$ as the embeddings of additional out-of-vocabulary special tokens, where one passage is represented as one special token, for example <p1> represents $p_{1}$.
@@
(1) $\mathbf{E}_{\text{In}}^{(1)}=\mathbf{E}_{t}(I\oplus q)\oplus\mathbf{E}_{M}(\bm{e}_{p_{1}})\cdots\oplus\mathbf{E}_{M}(\bm{e}_{p_{n}}),$
where $\mathbf{E}_{t}$ is the token embedding layer of LLM. The complete prompts are listed in Appendix A. In the next section, we will introduce how to output the ranking list in detail.
@@
3.2. Inference
During inference, listwise rerankers aim to output a ranking list directly. For LLM-based listwise approaches, we usually generate the ranking list autoregressively. In previous work, LLMs are prompted to generate a string that could be parsed into a ranking list, such as “[2] ¿ [3] ¿ [1]…” (Sun et al., 2023; Pradeep et al., 2023a). However, in early experiments, we found that generating a string representing ranking may be difficult and slow, as LLM may output in the wrong format or output useless content, such as explanation.
Input : Candidates $\mathcal{P}_{cand}=[p_{1},...,p_{n}]$, Initial Input Embeddings $\mathbf{E}_{\text{In}}^{(1)}$
@@ 3       $\bm{h}_{i}\leftarrow\text{Transformer}(\mathbf{E}_{\text{In}}^{(i)})$
Algorithm 1 Dynamic-Constrained Decoding
To address this issue, we propose a “Dynamic-Constrained Decoding” strategy in Algorithm 1. During decoding, we replace the original output layer over the whole vocabulary with the concatenation of embeddings of passages that need to be ranked, treating the embedding representation of those passages as a set of special tokens. Moreover, the decoding space, i.e., the output layer, is dynamically changed as the ranking process progresses, as fewer remaining passages need to be ranked, resulting in a continuous decrease in decoding space.
At each generation step, we no longer output a normal numerical token but instead constrain the decoding space only in these special tokens, to perform accurate ranking. Therefore, we can directly output a list of tokens that represent the ranking of passages, such as “<p2><p3><p1>…”. Furthermore, as the decoding space and the number of generated tokens are much smaller than the original vocabulary space, inference will be accelerated.
@@
It’s worth noting that there are existing works of constrained decoding (Willard and Louf, 2023), however, notable distinctions exist between our approach and theirs. Firstly, The decoding space of these related works is the original decoding space of LLM and is static, while that of our proposed method is outside the original vocabulary and dynamic. Secondly, These related works employed constrained decoding for generating text with strict format constraints. In contrast, our goal is simply to output a ranking list of tokens which leads to a more simple and more efficient method.
@@ 3.3. Training
(2) $\mathbf{E}_{\text{In-Align}}=\mathbf{E}_{t}(I)\oplus\mathbf{E}_{M}(\bm{e}_{t}),$
@@
(3) $\mathcal{L}_{\text{Align}}=-\sum_{i=1}\log P_{\theta}(t_{i}|\mathbf{E}_{\text{In-Align}}\oplus\mathbf{E}_{t}(t_{<i})).$
@@
Considering the decoding process, it can be viewed as a sequential ranking learning process: at each step, we provide the previously decoded rankings and maximize the probability of generating the next most relevant passage. Formally, if given a query $q$ and the golden ranking list $[p_{1},...,p_{n}]$, at step $i$, we maximize the conditional probability of $p_{i}$ given $q$ and $p_{<i}$:
(4) $P_{\theta}(p_{i}|q,p_{<i})$ $=P_{\theta}(p_{i}|\mathbf{E}_{\text{In}}^{(i)})$
@@
where $\theta$ is the model’s parameters. Considering the whole sequential process, it is equivalent to listwise learning to rank loss ListMLE (Xia et al., 2008):
(5) $\mathcal{L}_{\text{rank}}=-\sum_{i=1}^{n}\log P_{\theta}(p_{i}|\mathbf{E}_{\text{In}}^{(i)}).$
Here we only leverage the passage embeddings for ranking, as illustrated in the prompt (a) in Figure 3. The full prompts can be found in Appendix A.
Figure 3. Illustration of two types of training data and the learning-to-rank training process.
@@
(6) $\mathcal{L}_{\text{content}}=-\sum_{i=1}^{n}\log P_{\theta}(p_{i}|\mathbf{E}_{\text{In-c}}^{(i)}),$
@@
(7) $\mathcal{L}_{\text{KL}}=\sum_{i=1}^{n}{D}_{\text{KL}}(P_{\theta}(p_{i}|\mathbf{E}_{\text{In}}^{(i)})\|P_{\theta}(p_{i}|\mathbf{E}_{\text{In-c}}^{(i)}).$
@@
(8) $\mathcal{L}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{content}}+\alpha\mathcal{L}_{\text{KL}}.$
Here $\alpha$ is the hyperparameter. We fine-tune both MLP and LLM in this stage but keep the encoder frozen.
@@
4. Experiment Setup
4.1. Model selection